OpenStudy (anonymous):

If A+B+C=180。, prove that sin^2[A]+sin^2[b]+sin^2[C]=2+2cosAcosBcosC

OpenStudy (anonymous):

\(a +b+c =\pi\) \( b = \pi -(a+c)\) \[sin^2a + sin^2(\pi -(a+c) ) + sin^2c\] \[sin(a+c) = sina*cosc +cosa*sinc\] \[sin^2(a+c) = sin^2a*cos^2c + cos^2a*sin^2c + 2sina*cosa*cosc*sinc\] This should Help

OpenStudy (anonymous):

mellinia u must reply and give medal to Ishaan as he has atleast helped u

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

oh.